# The $\mathbb{S}_{n}$-equivariant Chow polynomial of the Braid matroid

### v2

@@ Abstract.
We determine the generating function for the $\mathbb{S}_{n}$-equivariant Chow polynomials of the braid matroid $B_{n}$, using a modular interpretation of this polynomial recently given by Devkota, Robotis, and Zahariuc. We also give a new modular interpretation of the numerical Chow polynomial of $B_{n}$ in terms of a certain moduli space of genus-zero relative stable maps to $\mathbb{P}^{1}$.
@@ 1. Introduction
$\{z_{i}=z_{j}\mid 1\leq i<j\leq n\}\subseteq\mathbb{C}^{n}.$
@@
Above, $\Lambda:=\mathbb{Q}[\![p_{1},p_{2},\ldots]\!]$ is the degree-completed ring of symmetric functions over $\mathbb{Q}$, and for a $\mathbb{S}_{n}$-representation $V$ we write $\operatorname{ch}_{n}(V)\in\Lambda$ for its Frobenius characteristic, which determines the decomposition of $V$ into irreducible $\mathbb{S}_{n}$-representations.
@@
$\mathcal{M}_{0,n+1}:=\frac{\mathrm{Conf}_{n+1}(\mathbb{P}^{1})}{\operatorname{Aut}(\mathbb{P}^{1})},$
@@
$\sum_{n\geq 0}h_{n}t^{n}=\exp\left(\sum_{n\geq 1}\frac{p_{n}}{n}t^{n}\right).$
@@
This is the usual Chow polynomial of the braid matroid, which has attracted significant attention in matroid theory: see §1.2 below.
@@ 55 $s_{5}(1+5t+5t^{2}+t^{3})+s_{41}(4t+4t^{2})+s_{32}(3t+3t^{2})+s_{221}(t+t^{2})$
In addition to the above formulas, we give a new modular interpretation of the Chow polynomial in terms of a certain moduli space of relative stable maps to $\mathbb{P}^{1}$.
@@
If $\mathcal{R}_{n}$ denotes the moduli space of genus-zero relative stable maps to $\mathbb{P}^{1}$ as in Definition 6.2, then there is an equality of dimensions
$\dim_{\mathbb{Q}}\operatorname{CH}^{k}(\mathcal{R}_{n})=\dim_{\mathbb{Q}}\operatorname{CH}^{k}(B_{n})$
for all $k\geq 0$.
Via Theorem C, the formulas in Corollary B apply to the moduli space $\mathcal{R}_{n}$, whose topology was previously studied by the first author in [Kan23].
@@ 1.1. Method
Our proof of Theorem A relies on a modular interpretation of the Chow polynomial $\mathrm{H}_{n}(t)$ as the $\mathbb{S}_{n}$-equivariant Chow polynomial of a certain moduli space $\mathcal{B}_{n}$ of projectivized multiscale differentials, which has recently been established by Devkota, Robotis, and Zahariuc [DRZ25]. We prove Theorem A by analyzing the classes of these moduli spaces in the Grothendieck ring of varieties. To consider all of the relevant moduli spaces at once, we use the formalism of $\mathbb{S}$-spaces developed by Getzler and Pandharipande [GP06]. To prove Theorem C, we prove that the classes of the moduli space $\mathcal{R}_{n}$ and $\mathcal{B}_{n}$ coincide in the Grothendieck ring of varieties, and conclude using the Chow-Künneth generation property as in [CL23].
@@ 1.2. Matroid-theoretic context
De Concini and Procesi constructed a so-called wonderful compactification of the complement of any complex arrangement of hyperplanes in [DCP95]. In a landmark paper, Feichtner and Yuzvinsky described the Chow ring of the wonderful compactification in terms of the underlying matroid [FY04]. This opened the door for defining the Chow ring $\operatorname{CH}^{\star}(M)$ of an arbitrary matroid $M$, even in the non-realizable case where it is not the Chow ring of some smooth projective variety. This Chow ring depends on the choice of a so-called building set for the lattice of flats. We employ the usual convention that $\operatorname{CH}^{\star}(M)$ is the Chow ring of $M$ with respect to the maximal building set.
@@
In a breakthrough article, Adiprasito, Huh and Katz established a Hodge theory for matroids by proving that $\operatorname{CH}^{\star}(M)$ satisfies the Kähler package, for any matroid $M$ [AHK18]. This lead to the resolution of the long-standing Heron–Rota–Welsh conjecture on the log-concavity of the coefficients of the characteristic polynomial of matroids. The development of the matroidal Hodge theory sparked several recent articles that study combinatorial questions on matroids via geometric tools, see e.g. [BHMPW22, ADH23].
@@
For a matroid $M$, we use the notation $\mathrm{H}^{\mathrm{num}}_{M}(t)$ for the Chow polynomial, which is the Hilbert–Poincaré series of $\operatorname{CH}^{\star}(M)$. The aforementioned work on the Hodge theory of matroids implies that $\mathrm{H}_{M}^{\mathrm{num}}(t)$ is a monic and palindromic polynomial with unimodal coefficients. The study of the Chow polynomial of a matroid has since gained significant attention. It is for instance conjectured that $\mathrm{H}_{M}^{\mathrm{num}}(t)$ has only real roots [FS24], and this was recently confirmed for all uniform matroids [BV25].
@@
Ferroni, Matherne, Stevens and Vecchi proved that the Chow polynomial is the inverse of the reduced characteristic polynomial in the incidence algebra of the lattice of flats [FMSV24]. This lead to a general recursion to compute the Chow polynomial of a matroid, previously obtained by Jensen–Kutler–Usatine [JKU21]. Their work also allowed for an even more general definition of Chow polynomials for posets [FMV24]. Using an $R$-labeling on the lattice of flats of a matroid, Stump produced a formula for the Chow polynomial of matroids using the descents along the chains in the lattice [Stu24]. Stump’s formula implies that the Chow polynomial is $\gamma$-positive.
@@
Recently, the authors of [EFMPV25] produced a recursive formula for the Hilbert–Poincaré series of the Chow ring of a matroid for an arbitrary building set. The Chow ring of the braid matroid of the minimal building set is the Chow ring of the moduli space $\overline{\mathcal{M}}_{0,n+1}$. They use their recursion to re-prove a formula for the Poincaré polynomial of the moduli space $\overline{\mathcal{M}}_{0,n+1}$ in terms of Stirling numbers of both kinds given by Aluffi, Marcolli, and Nascimento [AMN24].
@@ 1.3. Outline
In §2 we recall the moduli space of projectivized multiscale differentials as studied in [DRZ25]. In §3 we recall the formalism of $\mathbb{S}$-spaces developed by Getzler–Pandharipande, and reduce Theorem A to a combinatorial identity in the Grothendieck ring of $\mathbb{S}$-spaces. We then prove the identity in §4. We derive Corollary B in §5, and conclude by proving Theorem C in §6.
@@ 2. The moduli space of multiscale differentials
studied in [Dev24, DRZ25]. We first set up some notation for working with dual graphs of stable curves.
@@ 2.1. Stable trees
Given $\mathbf{T}\in\Gamma_{0,n}^{\star}$, define the canonical partial order $\leq_{\mathbf{T}}$ on the vertex set $V(\mathbf{T})$, by declaring $v_{1}\leq v_{2}$ if and only if there is a (possibly empty) directed path from $v_{1}$ to $v_{2}$ in $\mathbf{T}$. We write $v_{1}<_{\mathbf{T}}v_{2}$ if and only if there is a nonempty directed path from $v_{1}$ to $v_{2}$ in $\mathbf{T}$.
2.2. Moduli spaces
Let $\overline{\mathcal{M}}_{0,n+1}$ denote the moduli space of $(n+1)$-pointed stable curves of genus zero, where the marked points are labelled by the set $\{0,\ldots,n\}$. Each genus-zero, $(n+1)$-pointed stable curve $(C,p_{0},\ldots,p_{n})$ naturally has a dual tree $\mathbf{T}_{C}=(T_{C},m_{C})\in\Gamma_{0,n}^{\star}$; the vertex set $V(T_{C})$ is the set of irreducible components of $C$, the edge set $E(T_{C})$ is the set of nodes of $C$, and the marking function $m_{C}$ tracks the distribution of marked points among the irreducible components.
@@ 2.2. Moduli spaces
Let $(C,q_{0},\ldots,q_{n})$ be an $(n+1)$-pointed stable curve of genus zero, and let $C_{v}$ be an irreducible component of $C$, corresponding to a vertex $v\in V(\mathbf{T}_{C})$. Let $\eta$ be a node of $C$ which is contained in $C_{v}$. We say the node $\eta$ is incoming at $C_{v}$ if the edge $e\in E(\mathbf{T}_{C})$ representing $\eta$ is directed towards $v$, and we say $\eta$ is outgoing at $C_{v}$ otherwise. If $v$ is a non-root vertex of $\mathbf{T}_{C}$, then there is a unique node which is incoming at $C_{v}$, since $\mathbf{T}_{C}$ is a connected tree.
We can now define the space $\mathcal{B}_{n}$ from (2.1). The space $\mathcal{B}_{n}$ is a fine moduli space for tuples
@@
where
  • $(C,q_{0},\ldots,q_{n})$ is an $(n+1)$-pointed stable curve of genus 0 with marked dual tree $\mathbf{T}_{C}\in\Gamma_{0,n}^{\star}$;
  • $\ell:V(\mathbf{T}_{C})\to\{0,\ldots,k\}$ is a surjective map for some $k\geq 0$, such that whenever $v<_{T}w$, we have $\ell(v)<\ell(w)$;
  • if $v\in V(\mathbf{T}_{C})$ is the root vertex of $\mathbf{T}_{C}$ and $C_{v}$ denotes the corresponding irreducible component of $C$, then $\omega_{v}$ is a meromorphic differential on $C_{v}$ with a unique pole of order $2$ at $q_{0}$;
  • if $v\in V(\mathbf{T}_{C})$ is a non-root vertex and $C_{v}\subset C$ is the corresponding irreducible component, then $\omega_{v}$ is a meromorphic differential on $C_{v}$ with a unique pole of order $2$ at the incoming node of $C_{v}$.
The function $\ell$ is called a level structure on the dual graph $\mathbf{T}_{C}$, and we call the union of components $C_{v}\subset C$ with $\ell(v)=j$ the $j$th level of $C$. Each level of $C$ admits a diagonal $\mathbb{C}^{\star}$-action, which corresponds to rescaling the differential. Two tuples as in (2.6) are isomorphic if there is an isomorphism of the underlying pointed stable curves which is compatible with the choice of differentials, or if they differ by the level-by-level scaling of $\mathbb{C}^{\star}$. See [BCGGM24] for general background on moduli spaces of multiscale differentials, and [Dev24, DRZ25] for an in-depth study of the case of $\mathcal{B}_{n}$. The space $\mathcal{B}_{n}$ is a smooth projective variety of dimension n3n-3 [Dev24]. It admits an action of $\mathbb{S}_{n}$ which permutes the marked points $q_{1},\ldots,q_{n}$. The following theorem is implicit in the work of Devkota, Robotis, and Zahariuc [DRZ25].
@@
In fact, they prove that $\mathcal{B}_{n}$ is equal to the wonderful variety $Y_{n}$ determined by the maximal building set for $B_{n}$, by exhibiting both spaces as the same iterated blowup of $\overline{\mathcal{M}}_{0,n+1}$. The Chow ring of $B_{n}$ is by definition the Chow ring of $Y_{n}$ [FY04]. According to [DRZ25], Theorem 2.7 can also be deduced from the description of the log Chow ring of $\overline{\mathcal{M}}_{0,n+1}$ in [PRSS25, Theorem 11].

### v3

@@ Abstract.
We determine the generating function for the $\mathbb{S}_{n}$-equivariant Chow polynomials of the braid matroid $B_{n}$. The Chow polynomial of $B_{n}$ is the Poíncare polynomial of the wonderful compactification of the complement of the braid arrangement, with respect to the maximal building set. A key input to our result is the identification of this compactification with a certain moduli space of multiscale differentials, recently established by Devkota, Robotis, and Zahariuc. To prove our main formula, we study the classes of these moduli spaces in the Grothendieck ring of varieties via the formalism of $\mathbb{S}$-spaces developed by Getzler and Pandharipande. We also give a new interpretation of the numerical Chow polynomial of $B_{n}$ as the Poíncare polynomial of a moduli space of genus-zero relative stable maps to $\mathbb{P}^{1}$.
@@ 1. Introduction
(1.1) $\{z_{i}=z_{j}\mid 1\leq i<j\leq n\}\subseteq\mathbb{C}^{n}/\mathbb{C}\cdot(1,\ldots,1).$
@@
Above, $\Lambda:=\mathbb{Q}[\![p_{1},p_{2},\ldots]\!]$ is the degree-completed ring of symmetric functions over $\mathbb{Q}$ (see §3.2 for a precise definition), and for an $\mathbb{S}_{n}$-representation $V$ we write $\operatorname{ch}_{n}(V)\in\Lambda$ for its Frobenius characteristic. The element $\operatorname{ch}_{n}(V)$ determines the decomposition of $V$ into irreducible $\mathbb{S}_{n}$-representations.
@@
$\mathcal{M}_{0,n+1}:=\frac{\mathrm{Conf}_{n+1}(\mathbb{P}^{1})}{\operatorname{Aut}(\mathbb{P}^{1})}$
@@
$\sum_{n\geq 0}h_{n}t^{n}=\exp\left(\sum_{n\geq 1}\frac{p_{n}}{n}t^{n}\right),$
so e.g. $h_{1}=p_{1}$ and $h_{2}=(p_{1}^{2}+p_{2})/2$.
@@
This is the usual Chow polynomial of the braid matroid, which has attracted significant attention in matroid theory; see §1.2 below.
@@ 55 $s_{5}(1+5t+5t^{2}+t^{3})+s_{41}(4t+4t^{2})+s_{32}(3t+3t^{2})+s_{221}(t+t^{2})$
In addition to the above formulas, we give a new interpretation of the Chow polynomial as the Hilbert–Poíncare series of the Chow ring of a certain moduli space of relative stable maps to $\mathbb{P}^{1}$.
@@
If $\mathcal{R}_{n}$ denotes the moduli space of genus-zero relative stable maps to $\mathbb{P}^{1}$ as in Definition 6.2, then for all $k\geq 0$ there is an equality of dimensions
The Chow ring of $\mathcal{R}_{n}$ encodes the relative Gromov–Witten theory of $\mathbb{P}^{1}$ with maximal contact in the sense of [NR22]. Theorem C thus relates the combinatorics of the braid matroid with the enumerative geometry of $\mathbb{P}^{1}$. In particular, the formulas in Corollary B apply to the moduli space $\mathcal{R}_{n}$, whose topology was previously studied by the first author in [Kan23].
@@ 1.1. Method
Let $Y_{n}$ denote the wonderful compactification, in the sense of De Concini and Procesi [DP95], of the complement in $\mathbb{C}^{n}/\mathbb{C}\cdot(1,\ldots,1)$ of the braid arrangement (1.1), with respect to the maximal building set. Then
$\operatorname{CH}^{\star}(B_{n})=\operatorname{CH}^{\star}(Y_{n}).$
Our proof of Theorem A relies on a modular interpretation of $Y_{n}$ as a moduli space $\mathcal{B}_{n}$ of multiscale differentials in the sense of [BCGGM24]; the identification we use has recently been established by Devkota, Robotis, and Zahariuc [DRZ25]. We prove Theorem A by analyzing the classes of these moduli spaces in the Grothendieck ring of varieties. To consider all relevant moduli spaces at once, we use the formalism of $\mathbb{S}$-spaces developed by Getzler and Pandharipande [GP06]. To prove Theorem C, we prove that the classes of the moduli space $\mathcal{R}_{n}$ and $\mathcal{B}_{n}$ coincide in the Grothendieck ring of varieties, and conclude using the Chow-Künneth generation property as in [CL23].
@@ 1.2. Matroid-theoretic context
In a landmark paper, Feichtner and Yuzvinsky described the Chow ring of the De Concini–Procesi compactification of a hyperplane arrangement complement in terms of the underlying matroid [FY04]. Their work naturally led to the notion of the Chow ring $\operatorname{CH}^{\star}(M)$ of an arbitrary matroid $M$, even in the non-realizable case where it is not the Chow ring of some smooth projective variety. This Chow ring depends on the choice of a so-called building set for the lattice of flats. We employ the usual convention that $\operatorname{CH}^{\star}(M)$ is the Chow ring of $M$ with respect to the maximal building set.
@@
In a breakthrough article, Adiprasito, Huh and Katz established a Hodge theory for matroids by proving that $\operatorname{CH}^{\star}(M)$ satisfies the Kähler package, for any matroid $M$ [AHK18]. This led to the resolution of the long-standing Heron–Rota–Welsh conjecture on the log-concavity of the coefficients of the characteristic polynomial of matroids. Beyond this conjecture, a wealth of recent literature has shown that the matroidal Hodge theory is a robust geometric framework for studying a variety of combinatorial questions on matroids; see e.g. [BHMPW22, ADH23].
@@
For a matroid $M$, we use the notation $\mathrm{H}^{\mathrm{num}}_{M}(t)$ for the Chow polynomial, which is the Hilbert–Poincaré series of $\operatorname{CH}^{\star}(M)$. The aforementioned work on the Hodge theory of matroids implies that $\mathrm{H}_{M}^{\mathrm{num}}(t)$ is a monic and palindromic polynomial with unimodal coefficients. The study of the Chow polynomial of a matroid has gained significant attention. It is for instance conjectured that $\mathrm{H}_{M}^{\mathrm{num}}(t)$ has only real roots [FS24], and this was recently confirmed for all uniform matroids [BV25].
@@
Ferroni, Matherne, Stevens and Vecchi proved that the Chow polynomial is the inverse of the reduced characteristic polynomial in the incidence algebra of the lattice of flats [FMSV24]. This leads to a general recursion to compute the Chow polynomial of a matroid, previously obtained by Jensen–Kutler–Usatine [JKU21]. Their work also allowed for an even more general definition of Chow polynomials for posets [FMV24]. Using an $R$-labeling on the lattice of flats of a matroid, Stump produced a formula for the Chow polynomial of a matroid using the descents along the chains in the lattice [Stu25]. Stump’s formula implies that the Chow polynomial is $\gamma$-positive. The $\gamma$-positivity of the Chow polynomial is also proved in [FMSV24, Theorem 3.25].
@@
Recently, the authors of [EFMPV25] produced a recursive formula for the Hilbert–Poincaré series of the Chow ring of a matroid for an arbitrary building set. The Chow ring of the braid matroid with respect to the minimal building set is the Chow ring of the moduli space $\overline{\mathcal{M}}_{0,n+1}$. They use their recursion to re-prove a formula for the Poincaré polynomial of the moduli space $\overline{\mathcal{M}}_{0,n+1}$ in terms of Stirling numbers of both kinds given by Aluffi, Marcolli, and Nascimento [AMN25].
@@ 1.3. Outline
In §2 we recall the moduli space $\mathcal{B}_{n}$ of projectivized multiscale differentials as studied in [DRZ25], with a particular focus on the stratification of the space by so-called level trees. In §3 we recall $\mathbb{S}$-spaces and their basic properties established by Getzler–Pandharipande. We then reduce Theorem A to a combinatorial identity in the Grothendieck ring of $\mathbb{S}$-spaces. We prove this identity in §4, using the recursive structure of the level tree stratification. We derive Corollary B in §5 using basic properties of symmetric functions. Finally, we prove Theorem C in §6.
@@ 2. The moduli space of multiscale differentials
studied in [Dev26, DRZ25]. The moduli space $\mathcal{B}_{n}$ parameterizes stable $(n+1)$-marked curves of genus zero, together with certain choices of meromorphic differentials on the components of the curve. The choices of differentials are constrained by the dual tree of the stable curve. As such, we begin by setting some notation for working with dual trees.
@@ 2.1. Stable trees
Given $\mathbf{T}\in\Gamma_{0,n}^{\star}$, define the canonical partial order $\leq_{\mathbf{T}}$ on the vertex set $V(\mathbf{T})$, by declaring $v_{1}\leq v_{2}$ if and only if $v_{1}=v_{2}$ or there is a nonempty directed path from $v_{1}$ to $v_{2}$ in $\mathbf{T}$. We write $v_{1}<_{\mathbf{T}}v_{2}$ if and only if there is a nonempty directed path from $v_{1}$ to $v_{2}$ in $\mathbf{T}$.
2.2. The moduli space $\mathcal{B}_{n}$
Let $\overline{\mathcal{M}}_{0,n+1}$ denote the moduli space of $(n+1)$-pointed stable curves of genus zero, where the marked points are labelled by the set $\{0,\ldots,n\}$. Each genus-zero, $(n+1)$-pointed stable curve $(C,q_{0},\ldots,q_{n})$ naturally has a dual tree $\mathbf{T}_{C}=(T_{C},m_{C})\in\Gamma_{0,n}^{\star}$; the vertex set $V(T_{C})$ is the set of irreducible components of $C$, the edge set $E(T_{C})$ is the set of nodes of $C$, and the marking function $m_{C}$ tracks the distribution of marked points among the irreducible components.
@@ 2.2. The moduli space $\mathcal{B}_{n}$
Let $(C,q_{0},\ldots,q_{n})$ be an $(n+1)$-pointed stable curve of genus zero, and let $C_{v}$ be an irreducible component of $C$, corresponding to a vertex $v\in V(\mathbf{T}_{C})$. Let $\eta$ be a node of $C$ which is contained in $C_{v}$. We say the node $\eta$ is incoming at $C_{v}$ if the edge $e\in E(\mathbf{T}_{C})$ representing $\eta$ is directed towards $v$. If $v$ is a non-root vertex of $\mathbf{T}_{C}$, then there is a unique node which is incoming at $C_{v}$, since $\mathbf{T}_{C}$ is a connected tree.
We can now describe the space $\mathcal{B}_{n}$ from (2.1). We will only describe the closed points of the moduli space, since foundational aspects are treated in [BCGGM24]. The space $\mathcal{B}_{n}$ is a fine moduli space for tuples
@@
with the following properties.
  • $(C,q_{0},\ldots,q_{n})$ is an $(n+1)$-pointed stable curve of genus 0 with marked dual tree $\mathbf{T}_{C}\in\Gamma_{0,n}^{\star}$.
  • $\ell:V(\mathbf{T}_{C})\to\{0,\ldots,k\}$ is a surjective map for some $k\geq 0$, such that whenever $v<_{T}w$, we have $\ell(v)<\ell(w)$.
  • If $v\in V(\mathbf{T}_{C})$ is the root vertex of $\mathbf{T}_{C}$ and $C_{v}$ denotes the corresponding irreducible component of $C$, then $\omega_{v}\in H^{0}(C,\Omega_{C_{v}}(2q_{0}))$ is a non-zero meromorphic differential on $C_{v}$ with a unique pole of order $2$ at $q_{0}$.
  • If $v\in V(\mathbf{T}_{C})$ is a non-root vertex and $C_{v}\subset C$ is the corresponding irreducible component with incoming node $q_{v}\in C_{v}$, then $\omega_{v}\in H^{0}(C_{v},\Omega_{C_{v}}(2q_{v}))$ is a meromorphic differential on $C_{v}$ with a unique pole of order $2$ at $q_{v}$.
The function $\ell$ is called a level structure on the dual graph $\mathbf{T}_{C}$, and we call the union of components $C_{v}\subset C$ with $\ell(v)=j$ the $j$th level of $C$. Each level $j$ of $C$ admits a diagonal $\mathbb{C}^{\star}$-action, which corresponds to simultaneously rescaling the differentials $\{\omega_{v}\mid\ell(v)=j\}$. If the number of levels is $e$, this leads to a $(\mathbb{C}^{\star})^{e}$-action on such a tuple. Two tuples as in (2.6) are considered equivalent if they differ by the $(\mathbb{C}^{\star})^{e}$-action. This torus is usually called the level rotation torus in the literature on multiscale differentials.
Definition 2.7.
We refer to the tuples corresponding to closed points of $\mathcal{B}_{n}$ as in (2.6) as multiscale differentials. For such a tuple, the marked point $q_{0}$ is called the distinguished marking, and the marked points $q_{1},\ldots,q_{n}$ are the non-distinguished markings.
In [Dev26], Devkota shows that the space $\mathcal{B}_{n}$ is a smooth projective variety of dimension n3n-3. In [DRZ25, Theorem 3.1], Devkota, Robotis, and Zahariuc identify the moduli space $\mathcal{B}_{n}$ with the De Concini–Procesi compactification $Y_{n}$ of the complement of the braid arrangement (1.1) with respect to the maximal building set, by identifying both spaces with the same iterated blowup of $\overline{\mathcal{M}}_{0,n+1}$. Their identification immediately implies the following theorem.
@@
According to [DRZ25], Theorem 2.8 can also be deduced from the description of the log Chow ring of $\overline{\mathcal{M}}_{0,n+1}$ in [PRSS25, Theorem 11]. Theorem 2.8 lets us use the geometry of $\mathcal{B}_{n}$ to calculate the $S_{n}$-equivariant Chow polynomial of $B_{n}$.
2.3. Multiscale differentials on smooth curves
As established in [BCGGM24], the moduli space $\mathcal{B}_{n}$ is a simple normal crossings compactification of the subspace
$\mathcal{B}^{\circ}_{n}\subset\mathcal{B}_{n}$
where the underlying curve of the multiscale differential is smooth. Towards better understanding the geometry of $\mathcal{B}_{n}$, let us briefly describe $\mathcal{B}^{\circ}_{n}$. First, denote by
$\pi:\mathcal{C}_{0,n+1}\to\mathcal{M}_{0,n+1}$
the universal curve. Let $\omega_{\mathcal{C}_{0,n+1}/\mathcal{M}_{0,n+1}}$ denote the relative dualizing sheaf of this morphism. There are $n+1$ sections
$\mathrm{s}_{0},\ldots,\mathrm{s}_{n}:\mathcal{M}_{0,n+1}\to\mathcal{C}_{0,n+1}$
of $\pi$ determined by the marked points, and we can thus define the twist $\omega_{\mathcal{C}_{0,n+1}/\mathcal{M}_{0,n+1}}(2\mathrm{s}_{0})$. The pushforward
$\pi_{*}(\omega_{\mathcal{C}_{0,n+1}/\mathcal{M}_{0,n+1}}(2\mathrm{s}_{0}))$
is a line bundle on $\mathcal{M}_{0,n+1}$, with fiber over $(C,q_{0},\ldots,q_{n})$ given by $H^{0}(C,\Omega_{C}(2q_{0}))$: the higher cohomologies of the sheaves $\Omega_{C}(2q_{0})$ vanish, and $\dim_{\mathbb{C}}H^{0}(C,\Omega_{C}(2q_{0}))=1$ for all $C$.
Definition 2.9.
We write $\Omega\mathcal{M}_{0,n+1}$ for the total space of the line bundle $\pi_{*}(\omega_{\mathcal{C}_{0,n+1}/\mathcal{M}_{0,n+1}}(2\mathrm{s}_{0}))$ over $\mathcal{M}_{0,n+1}$. We also write $\Omega^{\star}\mathcal{M}_{0,n+1}$ for the complement of the zero section in $\Omega\mathcal{M}_{0,n+1}$.
The variety $\Omega^{\star}\mathcal{M}_{0,n+1}$ parameterizes irreducible smooth pointed genus-zero curves $(C,q_{0},\ldots,q_{n})$ together with a nonzero section of $H^{0}(C,\Omega_{C}(2q_{0}))$.
Lemma 2.10.
There are $\mathbb{S}_{n}$-equivariant isomorphisms
$\Omega^{\star}\mathcal{M}_{0,n+1}\cong\mathbb{C}^{\star}\times\mathcal{M}_{0,n+1}$
and
$\mathcal{B}^{\circ}_{n}\cong\mathcal{M}_{0,n+1}.$
Proof.
The Picard group $\mathrm{Pic}(\mathcal{M}_{0,n+1})$ is trivial, and $\Omega^{\star}\mathcal{M}_{0,n+1}$ is the complement of the zero section in the line bundle $\Omega\mathcal{M}_{0,n+1}$. Therefore $\Omega^{\star}\mathcal{M}_{0,n+1}\to\mathcal{M}_{0,n+1}$ is a trivial $\mathbb{C}^{\star}$-bundle, which is $\mathbb{S}_{n}$-equivariant by construction. The second isomorphism follows from the observation that
$\mathcal{B}^{\circ}_{n}\cong\Omega^{\star}\mathcal{M}_{0,n+1}/\mathbb{C}^{\star}$
where $\mathbb{C}^{\star}$ acts by scaling the differential. ∎
We will now describe the strata in the dual tree stratification of $\mathcal{B}_{n}$. The geometry of each stratum is concretely understood in terms of the trivial torus bundles $\Omega^{\star}\mathcal{M}_{0,n+1}$.